# Fast Polynomial Root-Finding and Approximation of the Eigenvalues of a Matrix

### v36

@@ 18 Bit operation complexity of general polynomial root-finding
For $q=d$ this is [81, Thm. 2], based on using novel data structure for hyperbolic approximation of $p$, proposed and analyzed by Guillaume Moroz in [81]. Trivially extend this theorem to the cases where $q>d$ (by considering $t(x)$ a polynomial of degree $q$ with $q-d$ leading coefficients 0) and where $q<d$ (by removing $d-q$ points $x_{i}$, with no increase of the complexity bound). ∎
We only apply this theorem for $\tau=0$, in which case its bit operation cost bound turns into
$\mathbb{B}_{\rm Moroz}=\tilde{O}((d+q)b)$ (18.1)
For comparison, the previous State of the Art was represented by the bound
$\mathbb{B}_{\rm Kirr}=\tilde{O}((d+q)(b+q)),$ (18.2)
due to [68, Thm. 3.9, Alg. 5.3, and Appendix A.3] by Peter Kirrinnis, which relied on combining his and Schönhage’s study of bit operation complexity of basic polynomial computations (cf. [134]) with the algorithm of [42, 84]. [^30] Precision of computing required in the latter algorithm grows fast as $d$ increases; e.g., the IEEE standard double precision of the customary numerical computations is not enough to support that algorithm already for $d>50$; for $b=O(\log(d))$ one can fix this deficiency by applying the Fast Multipole Method of Greengard and Rokhlin [50, 28, 105, 106] to the equivalent task of multiplication of Vandermonde matrix by a vector. One can, however, support bounds of Eqns. (18.1) and (18.2) on the overall bit operation complexity by computing with multiple precision for larger values $b$. Moroz obtained [81, Thm. 2] by applying bound (18.2) to his lower degree polynomials approximating $p(x)$.
The algorithms of [42, 84, 68, 81] extensively involve the coefficients of $p(x)$.
Bound (18.1) combined with the precision bound $O(\log(d))$ of Sec. 16 implies the bit operation cost bound $\tilde{O}((d+q)b)$ dominated by an estimated bit operation cost $\mathbb{B}$ for the evaluation of $p(x)$ with relative error $1/d^{O(1)}$ at $\tilde{O}(db)$ points. It remains to estimate $\mathbb{B}$, which is our next goal.
@@ 18.1 Reduction of our root-finders to multipoint polynomial approximation
In Thm. 18.1 a polynomial $t(x)$ with $||t(x)||_{1}\leq 1$ is approximated at some points $x\in D(0,1)$. To ensure the assumption about $x$, we scale the variable $x\mapsto Rx$ to map the discs $D(c_{\lambda},\rho_{\lambda}\sigma)\mapsto D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})\subseteq D(0,1)$, for $\bar{c}_{\lambda}=c_{\lambda}/R$, $\bar{\rho}_{\lambda}=\rho_{\lambda}/R\geq 1/2^{b}$, and all $\lambda$.
Then we write $t(x):=p(Rx)/\psi$ for $\psi=O(R^{d})$ such that $||t(x)||_{1}=1$ and approximate $t(x)$ within $1/2^{\bar{b}}$ or equivalently within the relative error bound $1/(|t(x)|2^{\bar{b}})$ at $\bar{m}$ points $x$ in $\bar{m}$ e/i tests. [^31] Under the Random Root Model of Sec. 1.5 we can whp decrease this bound by a factor of $\bar{m}/\log(\bar{m})$ and then extend this decrease to the related subsequent estimates. Due to Cor. 16.3 it is sufficient to ensure the bound $1/(|t(x)|2^{\bar{b}})=1/d^{O(1)}$; we achieve this by choosing $\bar{b}\geq\log(\frac{1}{|t(x)|})+O(\log(d))$.
In our e/i tests we only need to approximate $t(x)$ at the points $x$ lying on $\theta_{\lambda}$-isolated circles where $\frac{1}{\theta_{\lambda}-1}=O(m)$ (cf. Assumption 1 of Sec. 16.2). Then $\log(\frac{1}{|t(x)|})=O(d\log(m))$ by virtue of the bound of [134, Eqn. (9.4)]). [^32] That bound follows from the estimate $|f|\geq 1/2^{d}$ of [134, Thm. 4.2], which holds provided that $t(x)=f\prod_{j=1}^{m}(x-x_{j})\prod_{j=m+1}^{d}(\frac{x}{x_{j}}-1)$, $||t(x)||_{1}=1$, $|x_{j}|<1$ for $j\leq m$, and $|x_{j}|>1$ for $j>m$.
@@
For $m=d$, the above bound on $\log(\frac{1}{|t(x)|})$ is sharp up to a constant factor and is reached for the polynomial $t(x)=(x+1-\frac{1}{m})^{d}/(2-\frac{1}{m})^{d}$, $\rho_{\lambda}=1$, and $x=-1$. For $m<d$, [134, Thm. 4.5] implies a little stronger bounds, namely, $\log(\frac{1}{|t(x)|})=O(d+m\log(m))$ and hence $\bar{b}=O(m\log(m)+d+b)$ because all discs $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ lie in $\theta$-isolated unit disc $D(0,1)$ for $\theta-1$ exceeding a positive constant.
@@ 18.2 Bit operation cost of our e/i tests and root-finders
Apply Thm. 18.1 with $q$ and $b$ replaced by $\bar{q}=O(m^{2})$ and $\bar{b}=O(b+d+m\log(m))$, respectively, and obtain that the bit operation cost of our $O(m)$ e/i tests of Sec. 14 at any fixed subdivision step is in $\tilde{O}((b+d+m)m^{2})$ for $1\leq m\leq d$. This implies the upper estimate
@@
for the overall bit operation cost of our $O(b)$ subdivision steps. [^33] Under the Random Root Model of Sec. 1.5 we can decrease this bound by a factor of $m$ whp and can extend this decrease to the related subsequent estimates.
Based on [68, Alg. 5.3] and Eqn. (18.2) we obtain a little larger but still reasonably small upper bound

### v37

@@ 10.1 Straightforward extension of the precision bound and a direction to saving evaluation queries
For $q=d$ this is [64, Thm. 2], based on using novel data structure for hyperbolic approximation of $p$, proposed and analyzed by Moroz in [64]. Trivially extend this theorem to the cases where $q>d$ (by considering $t(x)$ a polynomial of degree $q$ with $q-d$ leading coefficients 0) and where $q<d$ (by removing $d-q$ points $x_{i}$, with no increase of the complexity bound). ∎
We only apply Thm. 10.1 for $\tau=0$, in which case its bit operation cost bound turns into
$\mathbb{B}_{\rm Moroz}=\tilde{O}((d+q)b).$ (10.2)
For comparison, the State of the Art before [64] was the bound
$\mathbb{B}_{\rm Kirr}=\tilde{O}((d+q)(b+q))$ (10.3)
of [52, Thm. 3.9, Alg. 5.3, and Appendix A.3] by Kirrinnis, who combined his and Schönhage’s study of the bit operation complexity of basic polynomial computations (cf. [106]) with the algorithm of [32, 67]. [^18] Precision of computing required in the latter algorithm grows fast as $d$ increases; e.g., the IEEE standard double precision of the customary numerical computations is not enough to support that algorithm already for $d>50$; for $b=O(\log(d))$ one can fix this deficiency by reducing the task to multiplication of Vandermonde matrix by a vector applying FMM (cf. [86, 87]). One can, however, support bounds of Eqns. (10.2) and (10.3) on the overall bit operation complexity by computing with multiple precision for larger values $b$. Moroz obtained [64, Thm. 2] by applying bound (10.3) to his lower degree polynomials approximating $p(x)$.
Bound (10.2) combined with the precision bound $O(\log(d))$ of Sec. 8 implies the bit operation cost bound $\tilde{O}((d+q)b)$, dominated by an estimated bit operation cost $\mathbb{B}$ for the evaluation of $p(x)$ with relative error $1/d^{O(1)}$ at $\tilde{O}(db)$ points. It remains to estimate $\mathbb{B}$, which we do next.
@@ 10.2 Reduction of our root-finders to multipoint polynomial approximation
In Thm. 10.1 a polynomial $t(x)$ with $||t(x)||_{1}\leq 1$ is approximated at some points $x\in D(0,1)$. To ensure the assumption about $x$, we scale the variable $x\mapsto Rx$ to map the discs $D(c_{\lambda},\rho_{\lambda}\sigma)\mapsto D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})\subseteq D(0,1)$, for $\bar{c}_{\lambda}=c_{\lambda}/R$, $\bar{\rho}_{\lambda}=\rho_{\lambda}/R\geq 1/2^{b}$, and all $\lambda$.
Then we write $t(x):=p(Rx)/\psi$ for $\psi=O(R^{d})$ such that $||t(x)||_{1}=1$ and approximate $t(x)$ within $1/2^{\bar{b}}$ or equivalently within the relative error bound $1/(|t(x)|2^{\bar{b}})$ at $\bar{m}$ points $x$ in $\bar{m}$ e/i tests. [^19] Under the Random Root Model of Sec. 1.5 we can whp decrease this bound by a factor of $\bar{m}/\log(\bar{m})$ and then extend this decrease to the related subsequent estimates. Due to Cor. 8.3 it is sufficient to ensure the bound $1/(|t(x)|2^{\bar{b}})=1/d^{O(1)}$; we deduce it by choosing $\bar{b}\geq\log(\frac{1}{|t(x)|})+O(\log(d))$.
In our e/i tests we only need to approximate $t(x)$ at the points $x$ lying on $\theta_{\lambda}$-isolated circles where $\frac{1}{\theta_{\lambda}-1}=O(m)$ (cf. Assumption 1 of Sec. 8.2). Then $\log(\frac{1}{|t(x)|})=O(d\log(m))$ by virtue of [106, Eqn. (9.4)]). [^20] That bound follows from the estimate $|f|\geq 1/2^{d}$ of [106, Thm. 4.2], which holds provided that $t(x)=f\prod_{j=1}^{m}(x-x_{j})\prod_{j=m+1}^{d}(\frac{x}{x_{j}}-1)$, $||t(x)||_{1}=1$, $|x_{j}|<1$ for $j\leq m$, and $|x_{j}|>1$ for $j>m$.
@@
For $m=d$, the above bound on $\log(\frac{1}{|t(x)|})$ is sharp up to a constant factor and is reached for the polynomial $t(x)=(x+1-\frac{1}{m})^{d}/(2-\frac{1}{m})^{d}$, $\rho_{\lambda}=1$, and $x=-1$. For $m<d$, [106, Thm. 4.5] implies a little stronger bounds, namely, $\log(\frac{1}{|t(x)|})=O(d+m\log(m))$ and hence $\bar{b}=O(m\log(m)+d+b)$ because all discs $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ lie in $\theta$-isolated unit disc $D(0,1)$ for $\theta-1$ exceeding a positive constant.
@@ 10.3 Bit operation cost of our e/i tests and root-finders
Apply Thm. 10.1 with $q$ and $b$ replaced by $\bar{q}=O(m^{2})$ and $\bar{b}=O(b+d+m\log(m))$, respectively, and obtain that the bit operation cost of our $O(m)$ e/i tests of Sec. 7 at any fixed subdivision step is in $\tilde{O}((b+d+m)m^{2})$ for $1\leq m\leq d$. This implies the upper estimate
@@
for the overall bit operation cost of our $O(b)$ subdivision steps. Based on [52, Alg. 5.3] and Eqn. (10.3) we obtain a little larger but still reasonably small upper bound